If 2 angles are supplements of each other and one angle is defined as 3r+12. What is its supplement angle?
Answer by ankor@dixie-net.com(22740)

If 2 angles are supplements of each other and one angle is defined as 3r+12.
What is its supplement angle?
:
"2 angles are supplements of each", just means the two angles add up to 180
:
Therefore the supplement angle (a) is:
a = 180 - (3r+12)
a = 180 - 3r - 12
a = 168 - 3r
,